# Interpretable Image Emotion Recognition: A Domain Adaptation Approach Using Facial Expressions

### v3

@@ 4.2 Training Strategy and Hyperparameter Setting
We have considered ‘happy,’ ‘anger,’ ‘hate,’ and ‘sad’ classes as these are the common emotion classes considered by the existing methods. The model training is performed using training-testing data split of $80$% and $20$%, $5$-fold cross-validation, a batch size of 16, a learning rate of $8\times 10^{-5}$ to $8\times 10^{-4}$ for up to $100$ epochs. The FER and IER models were first trained using cross-entropy loss and then trained together using the discrepancy loss, wher the Adam optimizer was used for both networks. Accuracy, defined as the proportion of true results (both true positives and true negatives) among the total number of cases examined, has been used as the evaluation metric. The emotion classes have been merged and re-labelled as per Plutchik’s emotion wheel [50]. The ‘contentment’ and ‘amusement’ labelled samples are re-labelled as ‘happiness,’ while the ‘disgust’ labelled samples are merged with the sample having a ‘hate’ label.
@@ 4.4.1 Accuracy and Confusion Matrices
The proposed IER system shows 60.98%, 58.86%, 69.13%, and 58.06% accuracies for IAPSa, ArtPhoto, FI, and EMOTIC datasets, respectively. Fig. 4 shows the corresponding confusion matrices.
@@ 4.4.2 Comparison With Existing Methods and Baselines
Table 1: Comparison of proposed system’s results with baselines and existing methods, across various datasets. The best and second-best results are highlighted in bold and underline, respectively.
@@
Baseline#1 42.51% 46.53% 57.66% 40.61%
Baseline#2 48.53% 50.26% 59.38% 44.54%
Baseline#3 47.93% 51.82% 62.81% 46.34%
Baseline#4 53.54% 49.96% 59.17% 48.89%
Baseline#5 57.37% 55.51% 62.77% 53.46%
@@
Multiple-instance Learning [51] 59.93% - - -
@@ 4.4.5 Ablation Studies
    Baseline#1 39.67% 40.17% 42.43% 42.36% 41.96%
    Baseline#2 47.53% 45.68% 45.17% 47.09% 48.53%
    Baseline#3 46.89% 46.26% 44.31% 47.16% 47.93%
    Baseline#4 53.54% 51.52% 50.29% 52.72% 49.48%
    Baseline#5 57.06% 52.29% 53.51% 57.37% 56.43%
    Proposed 56.14% 48.63% 51.36% 60.98% 58.93%
    ArtPhoto dataset.
@@
    Baseline#1 45.61% 43.51% 44.76% 46.07% 46.53%
    Baseline#2 50.17% 48.35% 49.47% 48.46% 50.26%
    Baseline#3 51.82% 49.96% 50.84% 50.83% 51.16%
    Baseline#4 47.37% 40.16% 45.46% 49.96% 42.43%
    Baseline#5 54.16% 52.71% 55.51% 54.37% 51.27%
    Proposed 50.53% 49.47% 58.86% 50.34% 56.96%
    FI dataset.
@@
    Baseline#1 56.72% 51.82% 52.17% 54.29% 57.66%
    Baseline#2 59.38% 56.46% 54.05% 58.86% 57.36%
    Baseline#3 58.24% 54.28% 50.11% 62.81% 56.32%
    Baseline#4 59.17% 58.74% 52.82% 57.61% 51.61%
    Baseline#5 61.62% 60.37% 62.75% 62.77% 61.53%
    Proposed 60.74% 54.47% 53.72% 67.17% 69.13%
    EMOTIC dataset.
@@
    Baseline#1 40.61% 39.53% 37.59% 40.54% 40.17%
    Baseline#2 44.54% 43.16% 42.36% 40.19% 43.96%
    Baseline#3 44.61% 45.46% 40.47% 42.93% 46.34%
    Baseline#4 46.89% 40.23% 42.87% 47.72% 41.86%
    Baseline#5 49.14% 50.37% 48.48% 52.53% 53.46%
    Proposed 53.16% 49.82% 52.97% 50.36% 58.06%
@@ 5 Conclusion and Future Work
This paper has introduced a novel feature-based domain adaptation strategy that effectively uses discrepancy loss to align the IER domain with the FER domain, employing the same network architecture for both tasks. Our approach not only demonstrates robust emotion recognition across diverse image contexts but also achieves notable accuracy metrics across multiple datasets. The integration of the DnCShap interpretability method enriches our understanding of key visual features influencing emotion recognition, enhancing transparency and trust in AI-driven systems. Looking ahead, the method’s potential applicability in areas lacking large-scale datasets invites further exploration, particularly in real-world scenarios where emotional contexts are highly variable. Future work should focus on refining the domain adaptation techniques to meet the nuanced demands of these applications and expanding the interpretability framework to include more complex emotional scenarios and diverse datasets. This ongoing research will aim to enhance the model’s scalability, adaptability, and efficiency, moving toward more sophisticated and universally applicable emotion recognition systems.
[…]
Acknowledgments The majority of this work, including conceptualization, implementation, result compilation, and initial manuscript preparation, was conducted at the Machine Vision Lab, Indian Institute of Technology Roorkee, India. The final stages of manuscript preparation and submission were carried out by the primary author while working at the University of Oulu, Finland.
@@
Conflicts of interest: Authors have no conflict of interest.

### v4

@@ 4.2.2 Training Strategy and Hyperparameter Setting
The model training is implemented in PyTorch, utilizing a training-testing data split of $80\%$ and $20\%$, $5$-fold cross-validation, with a standard batch size of 16, and a learning rate ranging from $8\times 10^{-5}$ to $8\times 10^{-4}$ for up to $100$ epochs. The FER and IER models were initially trained using cross-entropy loss and subsequently together using the discrepancy loss, with the Adam optimizer utilized for both networks. Accuracy, defined as the proportion of true results (both true positives and true negatives) among the total number of cases examined, has been used as the evaluation metric.
Hyperparameter Tuning: We employed Ray Tune [43] for hyperparameter optimization, allowing for a systematic and efficient search through the hyperparameter space. This approach facilitated the adjustment of learning rates dynamically, refinement of batch sizes to balance computational efficiency and model accuracy, and tuning of Adam optimizer parameters to optimize training dynamics. Different hyperparameter settings, such as beta values and epsilon for the Adam optimizer, were methodically varied to explore their impacts on the loss oscillations and the stability of model updates. Extended training durations beyond the standard $100$ epochs with early stopping based on validation loss were also tested to prevent overfitting while ensuring robust model performance. Inspired by other advanced hyperparameter optimization tools such as Optuna [2] and Keras Tuner [55], which are used extensively in various machine learning platforms, we incorporated similar strategies within our Ray Tune implementation to fine-tune the balance between cross-entropy and discrepancy loss. These modifications aided in refining the model’s capability to learn generalized features while effectively minimizing domain-specific discrepancies.
@@ 4.3 Models
  • Baseline#6: uses EfficientNet [74]. It is a scalable architecture that aims to achieve superior efficiency through compound scaling of depth, width, and resolution.
  • Baseline#7: It is based on the Vision Transformer (ViT) [16] model that applies attention mechanisms to segment the images into patches and processes them sequentially.
  • Baseline#8 uses Swin Transformer [46] which is a hierarchical transformer whose representation is computed with shifted windows, facilitating efficient modeling of image data at multiple scales to handle complex visual tasks.
@@ 4.4 Results
This section discusses the IER results along with the interpretable feature maps.
@@ 4.4.1 Accuracy and Confusion Matrices
The proposed IER system shows 61.86%, 62.47%, 70.78%, and 59.72% accuracies for IAPSa, ArtPhoto, FI, and EMOTIC datasets, respectively. Fig. 4 shows the corresponding confusion matrices.
@@ 4.4.2 Comparison With Existing Methods and Baselines
Table 2: Comparison of proposed system’s results with baselines and existing methods, across various datasets. The best and second-best results are highlighted in bold and underline, respectively. Standard deviations from five-fold cross-validation are shown in parentheses to indicate error margins.
@@
Baseline#1 42.83% (±1.5%) 47.37% (±1.2%) 58.21% (±1.8%) 41.17% (±2.0%)
Baseline#2 48.87% (±1.3%) 50.63% (±1.4%) 59.74% (±1.6%) 45.45% (±1.9%)
Baseline#3 48.31% (±1.6%) 52.26% (±1.1%) 63.17% (±1.3%) 46.57% (±1.7%)
Baseline#4 53.72% (±1.2%) 50.42% (±1.5%) 59.45% (±1.4%) 49.32% (±1.6%)
Baseline#5 57.46% (±1.0%) 55.87% (±1.3%) 63.12% (±1.1%) 54.37% (±2.1%)
Baseline#6 57.81% (±0.9%) 54.62% (±1.2%) 66.72% (±1.0%) 58.25% (±2.0%)
Baseline#7 59.13% (±1.1%) 55.24% (±1.4%) 65.86% (±1.2%) 57.65% (±1.8%)
Baseline#8 58.87% (±1.0%) 53.18% (±1.3%) 67.12% (±1.3%) 57.54% (±1.5%)
@@
Multi-instance Learning [60] 59.93% - - -
Multi-task DA [87] - 60.36% 63.92% -
CycleGAN [90] - 61.11% 63.87% -
CycleEmotionGAN [86] - 61.96% 67.78% -
Deep Cocktail Network [79] - 62.34% 65.31% -
Multi-source DA [44] - 63.58% 70.63% -
@@ 4.4.4 Layer-wise Interpretability
4.4.5 Computational Efficiency Analysis
Table 3 shows per-epoch computational times on an RTX 2080 GPU with a batch size of 16. We analyze four main operations in our pipeline: Baseline Training: The time to train the baseline model on each dataset for one epoch without domain adaptation. Domain Adaptation: The overhead introduced by our discrepancy-loss–based alignment between FER and IER, added at the end of each epoch. DnCShap: The additional cost for generating the Shapley-based attributions with Divide & Conquer Shap (DnCShap) each epoch. PCA Embeddings: Time to extract embeddings from the final layer and perform PCA for interpretability analyses.
Table 3: Per-epoch Computational Cost Analysis. Baseline training refers to the cost of plain supervised learning, while Domain Adaptation shows the extra overhead of our discrepancy-based alignment. Feature and Layer-wise Interpret reflect the per-epoch time to produce interpretability outputs. Increase percentages compared to the baseline are shown with an upward arrow.
Component Operation IAPSa ArtPhoto FI EMOTIC
Baseline Training Training Time (s) 4 6 25 28
Domain Adaptation Adaptation Time (s) 1.2 1.9 7.2 9.3
Feature-wise Interpret Analysis Time (s) 0.5 0.6 2.8 3.2
Layer-wise Interpret Analysis Time (s) 0.3 0.4 1.7 1.9
[…]
We observe that smaller datasets (IAPSa and ArtPhoto) require only a few seconds per epoch, whereas larger datasets (FI and EMOTIC, each containing over 20,000 images) demand correspondingly more time. Despite the discrepancy-loss domain adaptation overhead, it remains only a modest fraction ($\sim$20–40%) of baseline training. DnCShap analysis takes a few seconds at most, owing to its linear-time complexity. PCA computations are also minimal. Overall, even with additional domain adaptation and interpretability steps, our method remains computationally tractable for practical use, completing training in a reasonable time (on the order of minutes per epoch) for the smaller datasets and under a minute per epoch for the larger FI and EMOTIC datasets.
4.4.6 Human Annotation Study
To evaluate how well our model’s interpretations align with human perception of non-facial or generic images, we conducted a small-scale annotation study on a subset of each dataset (IAPSa, ArtPhoto, FI, and EMOTIC). We randomly selected 100 images per dataset that primarily depict scenes, objects, or contexts without clear human faces.
Annotation Procedure: Three volunteers (postgraduate students with familiarity in affective computing) independently labeled each image with one of the four emotions (happy, sad, hate, anger) and indicated the visual cues driving their label choice. They received no information about model outputs.
Inter-Annotator Agreement: We used Cohen’s Kappa ($\kappa$) [76] to quantify pairwise agreement. Table 4 shows the average $\kappa$ across annotator pairs. The moderate agreement for FI and EMOTIC underscores the subjective nature of emotional cues in real-world or context-heavy scenes.
Comparison with Model: We compared the model’s predicted labels to the majority vote among annotators for each image. In 78–83% of cases, the model matched the humans’ dominant label. Additionally, DnCShap-highlighted regions overlapped with human-indicated cues (e.g., color or foreground objects) in 70% of instances.
Table 4: Human Annotation Study on 50 Non-Facial Images per Dataset. Model-Human Align is the proportion of images where the model prediction matched the majority human label.
Dataset Avg. Cohen’s ($\kappa$) Model-Human Align Annotators’ Agreement
IAPSa 0.79 81.16% Moderate-High
ArtPhoto 0.76 78.65% Moderate
FI 0.81 83.23% High
EMOTIC 0.75 79.64% Moderate
Overall, this study indicates that our model’s predictions and saliency outputs provide a reasonable correspondence with human judgments in generic, non-facial images. However, the moderate agreement levels among annotators also highlight the inherent subjectivity in emotional perception. Future extensions could expand the subset size and include a broader pool of annotators or alternative emotion categories.
@@ 4.4.7 Ablation Studies
    Baseline#1 40.24% 40.23% 42.43% 42.83% 42.24%
    Baseline#2 48.27% 46.24% 45.17% 48.87% 49.25%
    Baseline#3 47.45% 46.53% 44.31% 48.31% 48.43%
    Baseline#4 53.73% 51.63% 50.29% 53.72% 50.25%
    Baseline#5 57.16% 52.57% 53.51% 57.46% 56.54%
    Baseline#6 56.62% 51.34% 52.93% 56.72% 57.81%
    Baseline#7 55.34% 51.67% 51.47% 59.13% 58.41%
    Baseline#8 56.72% 52.34% 53.24% 57.93% 58.87%
    Proposed 57.24% 53.13% 53.36% 61.86% 58.93%
    ArtPhoto dataset
@@
    Baseline#1 45.87% 47.73% 46.62% 46.28% 46.75%
    Baseline#2 50.35% 50.85% 51.75% 48.75% 50.24%
    Baseline#3 52.13% 52.37% 52.42% 51.18% 51.34%
    Baseline#4 47.82% 51.14% 48.68% 50.23% 42.62%
    Baseline#5 54.16% 56.16% 57.16% 54.75% 51.45%
    Baseline#6 53.34% 55.74% 58.22% 53.28% 52.34%
    Baseline#7 54.56% 56.12% 60.14% 54.62% 58.42%
    Baseline#8 53.82% 55.89% 59.83% 55.34% 57.29%
    Proposed 54.62% 56.47% 62.47% 56.34% 60.27%
    FI dataset
@@
    Baseline#1 56.37% 51.97% 52.63% 54.55% 58.21%
    Baseline#2 59.52% 56.56% 54.56% 62.83% 59.74%
    Baseline#3 58.17% 55.12% 51.07% 63.18% 63.17%
    Baseline#4 59.28% 56.97% 53.18% 57.86% 59.45%
    Baseline#5 60.56% 56.64% 54.35% 63.13% 63.12%
    Baseline#6 59.96% 57.62% 55.67% 64.53% 66.72%
    Baseline#7 60.12% 56.94% 56.35% 65.12% 65.86%
    Baseline#8 60.43% 57.12% 55.98% 64.94% 67.12%
    Proposed 61.53% 57.67% 56.72% 68.17% 70.78%
    EMOTIC dataset
@@
    Baseline#1 40.85% 39.83% 38.25% 40.93% 41.17%
    Baseline#2 44.94% 43.63% 42.68% 42.93% 45.45%
    Baseline#3 44.71% 46.73% 41.29% 43.64% 46.32%
    Baseline#4 46.99% 40.83% 43.73% 50.87% 47.53%
    Baseline#5 49.44% 50.73% 49.18% 52.86% 54.37%
    Baseline#6 51.46% 52.64% 51.64% 53.75% 58.25%
    Baseline#7 51.74% 53.64% 50.86% 55.65% 57.94%
    Baseline#8 50.96% 51.86% 51.12% 55.54% 57.83%
    Proposed 53.16% 54.29% 53.33% 56.36% 59.72%
@@ 4.5 Discussion
Generalizability Across Datasets: Our proposed method has demonstrated robust performance across standard datasets used in facial emotion recognition. However, the scalability and generalizability of the system to other datasets, especially those that might contain more complex or abstract emotional cues, warrant further exploration. The architecture and methodologies employed in our system are designed to be adaptable and flexible, allowing for potential extensions to datasets encompassing a broader range of emotional expressions and cultural variations. For example, the inclusion of domain adaptation techniques such as discrepancy loss ensures that our model can learn domain-invariant features, which is crucial for applications involving diverse emotional datasets. Future work will focus on testing the model’s performance on datasets like the Emotionet, which includes a wide array of subtle and complex emotional expressions, often captured in less controlled environments. Additionally, we plan to explore the integration of multimodal data sources, such as audio and textual information, to enrich the emotional context and enhance the model’s interpretability and adaptability to real-world scenarios where emotional cues are not solely visual.
Failure Case Analysis: Despite achieving competitive accuracy across multiple datasets, our model exhibits certain failure cases that warrant further analysis. A key challenge arises in the misclassification of visually similar emotions, such as confusion between fear and surprise or sadness and neutral expressions, particularly in the FER domain. In the IER setting, errors frequently occur in images where emotional cues are highly context-dependent rather than explicitly conveyed through facial expressions. Additionally, domain adaptation limitations contribute to errors when transferring knowledge from FER to IER, especially when the dataset distributions differ significantly. Low-quality images, occlusions, and dataset biases further affect the model’s performance, leading to occasional inconsistencies in predictions. To mitigate these issues, potential improvements include incorporating uncertainty-aware learning, improving dataset balance, and leveraging multi-modal cues beyond visual information to enhance emotion recognition reliability. This analysis highlights the importance of refining adaptation strategies and dataset curation to address these failure cases effectively.
@@ 5 Conclusion and Future Work
This paper has introduced a novel feature-based domain adaptation strategy that effectively uses discrepancy loss to align the IER domain with the FER domain, employing the same network architecture for both tasks. Our approach not only demonstrates robust emotion recognition across diverse image contexts but also achieves notable accuracy metrics across multiple datasets. The integration of the DnCShap interpretability method enriches our understanding of key visual features influencing emotion recognition, enhancing transparency and trust in AI-driven systems. Looking ahead, the method’s potential applicability in areas lacking large-scale datasets invites further exploration, particularly in real-world scenarios where emotional contexts are highly variable. Future work should focus on refining the domain adaptation techniques to meet the nuanced demands of these applications and expanding the interpretability framework to include more complex emotional scenarios and diverse datasets. Additionally, exploring the integration of multimodal data, such as audio and textual cues, could provide a more comprehensive understanding of emotional states and further improve the accuracy and robustness of emotion recognition systems. This ongoing research will aim to enhance the model’s scalability, adaptability, and efficiency, moving toward more sophisticated and universally applicable emotion recognition systems.
@@
Acknowledgments: The majority of this work was conducted at the Machine Vision Lab, Indian Institute of Technology Roorkee, India. The final stages of manuscript preparation and submission were carried out by the primary author while working at the University of Oulu, Finland.
@@
Competing interests: Authors have no conflict of interest.
@@
Appendix A
The proposed interpretability DnCShap has been mathematically illustrated in Algorithm 1, which uses Algorithm 2 for recursive calculations.
Define $model$: DNN model,
Define $data$: Image pixels,
Define $wd,ht$: Image’s width & height,
Output $SHAP\_value$.
procedure DnCShap ($model,\ data,\ wd,\ ht,\ times$)
$data\_b=np.zeros([wd,ht,3])$
$data\_f=data$
$data\_f=data\_f.reSHAPe(1,wd,ht,3)$
$data\_b=data\_b.reSHAPe(1,wd,ht,3)$
$\triangleright$ Perturb by dividing image into two parts, along x or y axis
$data\_1,data\_2=get\_parts(data,width,height,x)$
$pred=model.predict(data\_f)$
$arg\_{max}=np.argmax(pred)$
$pred\_f=pred[0][arg\_max]$
$pred\_b=model.predict(data\_b)[0][arg\_max]$
$pred\_1=model.predict(data\_1)[0][arg\_max]$
$pred\_2=model.predict(data\_2)[0][arg\_max]$
$score\_1=((pred\_1-pred\_b)+(pred\_f-pred\_2))/2$
$score\_2=((pred\_2-pred\_b)+(pred\_f-pred\_1))/2$
$SHAP\_value=np.zeros([width,height])$
$times=times-1$
RecSHAP ($model,data,wd,ht,times,1,0,x\_m,0,ht,pred\_b+pred\_2,$
$pred\_1+pred\_f,score\_1,arg\_max,SHAP\_value$)
RecSHAP ($model,data,wd,ht,times,1,x\_m,wd,0,ht,pred\_b+pred\_1,$
$pred\_2+pred\_f,score\_2,arg\_max,SHAP\_value$)
return $SHAP\_value$
Algorithm 1 DnCShap
Define $model$: DNN model,
Define $data$: Image pixels,
Define $wd,ht$: Image’s width & height,
Define $pred\_b,pred\_f$: prediction with both/no parts perturbed,
Define $x\_s,x\_e$ initial & final dimensions for width
Define $y\_s,y\_e$ initial & final dimensions for height
Define $times$: number of parts the image is to be divided,
Define $arg\_{max}$: predicted label,
Define $score$: SHAP value of the part,
Define $value$: activation map.
procedure RecSHAP ($model,\ data,\ wd,\ ht,\ times,\ i,x\_s,\ x\_e,$
$y\_s,\ y\_e,\ pred\_b,\ pred\_f,\ score,\ arg\_{max},\ value$)
if (times == 0) then
    $value[x_{s}:x_{e},y_{s}:y_{e}]$ = score;
else (if i == 1) then
    z = y;
else
    z = x;
$\triangleright$ Perturb by dividing image into two parts, along x or y axis
$data\_1,data\_2=get\_parts(data,\ wd,\ ht,\ z)$
$pred\_1=model.predict(data\_1)[0][arg\_{max}]$
$pred\_2=model.predict(data\_2)[0][arg\_{max}]$
$score\_1=(((pred\_1-pred\_b)+(pred\_f-pred\_2))/2)/2$
$score\_2=(((pred\_2-pred\_b)+(pred\_f-pred\_2))/2)/2$
$times=times-1$
$i=abs(i-1)$
RecSHAP ($model,\ data,\ wd,\ ht,\ times,\ i,\ x\_s,\ x\_e,\ y\_s,\ y\_e$,
$pred\_b+pred\_2,\ pred\_1+pred\_f,\ score\_2,\ arg\_{max},\ value$)
RecSHAP ($model,\ data,\ wd,\ ht,\ times,\ i,\ x\_s,\ x\_e,\ y\_s,\ y\_e$,
$pred\_b+pred\_1,\ pred\_2+pred\_f,\ score\_2,\ arg\_{max},\ value$)
Algorithm 2 RecSHAP (recursive procedure)